OpenStudy (harrisonshark):

The endpoints of the diameter of a circle are (-7, 3) and (5, 1). What is the center of the circle?

OpenStudy (anonymous):

\[midpoint=\frac{ x _{1}+x _{2} }{ 2 }, \frac{ y _{1}+y _{2} }{ 2 }\] M=-7+5/2, 2+1/2 m=-2/2, 4/2 m=-1,2. the midpoint or centre is at (-1,2)